# Disentangling Core and Edge Mechanisms of the L-mode Density Limit in DIII-D

### v2

@@ III.3.1 Development of Radiation Front at the HFS edge
Figure 7 shows the temporal evolution of the radiated power distribution, reconstructed from bolometer measurements in a high-power discharge. This radiative feature is geometrically similar to a MARFE (Multifaceted Asymmetric Radiation From the Edge) [16, 36]. However, in these experiments, it manifests as a non-disruptive, quasi-stable radiation front that leads to enhanced fluctuations and density saturation. The intense but localized radiative edge cooling has also been observed in divertor detachment experiments in NT discharges [37].
Initially (Fig. 7(a), $t=2800$ ms), before the density ramp-up, radiation is concentrated near the inner divertor target, typical of an attached divertor regime. As the density is raised (Fig. 7(b), $t=3800$ ms), enhanced radiation emerges near the X-point, signaling the development of the radiative front. With a further density increase (Fig. 7(c), $t=4800$ ms), the condensation front becomes fully developed and moves upward along the high-field side separatrix, a process strongly correlated with the edge cooling shown previously in Fig. 6. Finally, the radiation front propagates further towards the HFS (inboard), and the radiated power near the separatrix intensifies (Fig. 7(d), $t=5800$ ms). It is noteworthy that in this discharge the disruption occurs shortly after a drop in the NBI power (at $t\approx 5800$ ms), which triggered an $n=1$ locked mode, and is not directly caused by the radiative edge cooling.
III.3.2 Edge Density Saturation by Radiative Instabilities
Previous observations suggest that the edge density saturates as radiative cooling intensifies. To quantify this relationship, we compare the evolution of the edge density with the amplitude of line-integrated density fluctuations. As the interferometer chord on the high-field side (HFS) passes directly through the developing radiation front, its signal serves as a reliable proxy for the intensity of the radiative instability.
@@ III.3.2 Edge Density Saturation by Radiative Instabilities
Figure 8 plots the edge density (at $\rho=0.95$) against the RMS amplitude of density fluctuations on a log-log scale. The data are collected from multiple discharges and colored by the energy confinement factor $H_{98y2}$, reveal a sharp bifurcation between two distinct operational regimes.
In the low-turbulence regime, characterized by high $H_{98y2}$ values, the edge density increases nearly linearly with HFS turbulence, following a power law of $n_{e,\text{edge}}\propto\tilde{n}_{\text{HFS}}^{1.01}$ (gray line in Fig. 8(a)). This indicates a phase where the increasing density turbulence does not yet limit the achievable edge density and plasma confinement.
This relationship breaks down abruptly as the HFS turbulence exceeds a critical threshold, marking a transition to a high-turbulence regime. Here, the fluctuation amplitude, now dominated by the radiative instability itself, continues to grow by nearly an order of magnitude. In stark contrast, the edge density saturates completely, exhibiting a nearly flat scaling of $n_{e,\text{edge}}\propto\tilde{n}_{\text{HFS}}^{0.1}$ (cyan line in Fig. 8(b)). Critically, this density saturation is accompanied by a sharp degradation in confinement, with $H_{98y2}$ dropping significantly. The overall scaling relationship is $n_{e,\text{edge}}\propto\tilde{n}_{\text{HFS}}^{0.26}$. This analysis suggests that the edge density limit is not a gradual process, but a sharp transition triggered by enhanced turbulence, likely driven by a radiative instability and the associated edge cooling. This instability clamps the edge density at a level consistently below the Greenwald limit and can be related to the loss of the edge confinement.
III.3.3 Turbulence Dynamics and the Breakdown of Zonal Flow Regulation
While the preceding analysis points to the radiative instability as the primary limit on edge density, it is crucial to understand the interplay between the turbulence and the developing radiation front. For this purpose, we use 2D measurements from the BES diagnostic, which provides a detailed look at ion-scale density fluctuations. This analysis utilizes a lower-power discharge, where the radiative cooling occurs at lower density, allowing for higher-quality BES signals.
@@ III.3.3 Turbulence Dynamics and the Breakdown of Zonal Flow Regulation
Figure 9 shows the evolution of key turbulence characteristics at the LFS edge as the radiative front forms (onset around 3700 ms, dashed purple curves). A fundamental change in the turbulence character is signaled by the dimensionless adiabaticity parameter [38, 13], $\alpha_{\text{adia}}=k_{\parallel}^{2}v_{te}^{2}/(\omega\nu_{ei})$. Here, $k_{\parallel}=1/qR$ is the parallel wavenumber, $v_{te}$ is the electron thermal velocity, $\nu_{ei}=\frac{\sqrt{2}e^{4}n_{i}\ln\Lambda}{12\pi^{3/2}\epsilon_{0}^{2}m_{e}^{1/2}T_{e}^{3/2}}$ is the electron-ion collision frequency, and $\omega$ is the turbulence frequency. As the edge cools, $\alpha_{\text{adia}}$ drops below unity, indicating a transition to a non-adiabatic regime, which is known to enhance turbulence and weaken zonal flow generation [13, 14].
Consistent with this transition, the normalized density fluctuation amplitude ($\tilde{n}/\bar{n}$, Fig. 9(b)) grows steadily. The mean poloidal phase velocity ($\bar{V}_{\theta}$, Fig. 9(c)) remains relatively unchanged until the onset of radiative cooling. Crucially, the ability of the turbulence to self-regulate via zonal flows collapses. This is evident in the Reynolds stress ($\langle\tilde{v}_{\theta}\tilde{v}_{r}\rangle$, Fig. 9(d)), the primary drive for zonal flows. The Reynolds stress flattens, indicating that the nonlinear transfer of energy from micro-turbulence to stabilizing macroscopic flows, $\bar{V}_{\theta}\partial_{r}\langle\tilde{v}_{\theta}\tilde{v}_{r}\rangle$, has been severely compromised.
@@
The consequence of this weakly regulated turbulence is an increase in outward transport. The turbulent particle flux ($\langle\tilde{n}\tilde{v}_{r}\rangle$, Fig. 9(e)) is directed outward and grows, while the turbulent spreading flux ($\langle\tilde{n}^{2}\tilde{v}_{r}\rangle$, Fig. 9(f)) indicates that turbulence intensity is spreading inward from the separatrix. These detailed measurements indicate a potential causal sequence: edge cooling drives a transition to a non-adiabatic turbulent state where the zonal flow self-regulation mechanism fails. This leads to enhanced outward transport, providing the micro-physical link between the onset of strong radiative cooling and the ultimate saturation of the edge density.
III.4 The Core Density Limit: Role of Avalanche-like Turbulence
While the edge radiative condensation directly limits the edge density, the core density continues to increase, even surpassing the Greenwald density after the radiative front has formed. This divergence strongly suggests that a different, core-localized mechanism is at play. We now present evidence that this mechanism is a fundamental change in the nature of core turbulent transport.
III.4.1 Signatures of Avalanche-like Transport in the Core
The collapse of stabilizing $E_{r}\times B$ shear (as shown in Fig. 5) allows enhanced core turbulence. This is confirmed by examining line-integrated density fluctuations from the RIP diagnostic at the midplane. As the plasma approaches its density limit, the statistical character of this turbulence undergoes a significant change.
@@ III.4.1 Signatures of Avalanche-like Transport in the Core
Figure 10(a) shows the evolution of the probability density function (PDF) of these fluctuation amplitudes (high-pass filtered at $f>5$ kHz to avoid machine vibration noise). Early in the density ramp-up, the PDF is Gaussian (yellow), characteristic of diffusive, random-walk-like transport. However, as the density limit is approached—and particularly after the onset of the edge radiative cooling—the PDF develops pronounced “heavy tails” (purple). This indicates a transition to a state dominated by large-amplitude, intermittent transport events, a clear departure from simple diffusive behavior.
This transition is quantified by the kurtosis of the PDF (Fig. 10(b)), a measure of its “tailedness.” The kurtosis rises significantly after the onset of the radiative cooling and continues to increase, confirming the growing intermittency. Further evidence comes from the power spectrum of the fluctuations (Fig. 11(a)). After the radiative cooling onset, a distinctive “$1/f$-type” power-law scaling emerges at low frequencies. This type of spectrum is a recognized signature of self-organized criticality (SOC) and avalanche-like dynamics [39, 40, 41], where small perturbations can trigger cascading transport events of all sizes.
@@
Finally, the transition to an avalanche-like state is supported by the Hurst exponent ($H$), a measure of long-range temporal correlations in the data (Fig. 11(b)). The Hurst exponent rises from $H\approx 0.5$ (indicative of uncorrelated, random events) to $H\approx 0.8$ (indicative of persistent, correlated dynamics) as the density limit is approached. Consistently, the energy confinement factor $H_{98y2}$ also decreases substantially when the Hurst exponent $H$ begins to increase, indicating a loss of confinement associated with this transition. Collectively, the heavy-tailed PDFs, increased kurtosis, $1/f$ spectrum, and elevated Hurst exponent point to a consistent picture: core transport transitions from a diffusive to an intermittent, avalanche-like regime, which is ultimately responsible for limiting the core density.
@@ III.4.2 Core Density Saturation by Avalanche-like Turbulence
The evidence for a transition to avalanche-like transport leads to a clear hypothesis: this enhanced turbulence is the primary factor limiting the core density. To test this hypothesis, we examine the direct relationship between the core density (at $\rho\approx 0.7$) and the amplitude of the midplane turbulence, as shown in Fig. 12. The data from multiple discharges, plotted on a log-log scale and colored by the energy confinement factor ($H_{98y2}$), reveal a distinct bifurcation between two transport regimes.
In the low-turbulence regime, corresponding to high confinement ($H_{98y2}>1.0$, yellow data points), the core density rises robustly with the turbulence amplitude, following a relatively strong power-law scaling of $n_{\text{e,core}}\propto\tilde{n}_{rms}^{0.49}$. This represents a phase where the core can still be effectively fueled, and the density increases despite the presence of turbulence.
@@
However, as the avalanche-like transport intensifies and confinement degrades ($H_{98y2}<1.0$, purple data points), this relationship changes dramatically. The scaling becomes substantially weaker, flattening to $n_{\text{e,core}}\propto\tilde{n}_{\text{rms}}^{0.17}$. This transition to a weaker scaling indicates that the avalanche-like density fluctuations increase turbulent transport so effectively that they prevent significant increases in the core density. Notably, this saturation occurs at densities well above the Greenwald limit, reaching nearly twice the Greenwald density.
The saturation of the core density is more clearly visible in the linear-scale plot in Fig. 12(b). The overall response of the core density to the turbulence amplitude scales as $n_{\text{core}}\propto\tilde{n}_{\text{rms}}^{0.26}$, which also underscores the saturated behavior with increasing turbulence. The weak dependence on turbulence in this regime aligns well with the macroscopic power scaling of the core density limit ($n_{\mathrm{e,core}}\propto P_{\mathrm{SOL}}^{0.27\pm 0.03}$) found earlier. This alignment, combined with the avalanche-like dynamics, suggests that the core density limit is governed by a fundamental transition in the core turbulent transport itself, which ultimately saturates the density profile.
@@ IV Conclusion and Discussion
This study has re-examined the L-mode density limit by leveraging the unique properties of negative triangularity plasmas on the DIII-D tokamak. Our findings suggest that the density limit, at least in this configuration, is not a singular disruptive boundary, but a multifaceted saturation phenomenon governed by distinct physical mechanisms: an edge density limit due to radiative edge cooling, and a core density limit driven by avalanche-like turbulent transport. Moreover, this study demonstrated sustained, non-disruptive operation at densities up to 1.8 times the Greenwald limit with substantial auxiliary heating.
At the plasma edge, the density limit manifests as an abrupt saturation event. This limit is triggered by the onset of enhanced turbulence driven by a radiative instability on the high-field side. While geometrically similar to a disruptive MARFE, this radiative instability acts as a quasi-stable front that clamps the edge density at a level below the Greenwald limit, without terminating the discharge.
In contrast, the core density is not directly constrained by the edge cooling and continues to rise well beyond the Greenwald limit. Its saturation is instead governed by a fundamental shift in the nature of core turbulence. As the density increases, core transport transitions from a diffusive state to one characterized by intermittent, avalanche-like dynamics. This is supported by signatures including heavy-tailed PDFs, $1/f$-type power spectra, and elevated Hurst exponents. This highly persistent, non-diffusive transport regime ultimately prevents further increases in the core density.
@@
Furthermore, these results suggest that high Greenwald fraction operation is feasible in future tokamaks in negative triangularity configurations, provided that the edge cooling and core turbulence dynamics are well understood and controlled. Specifically, innovative divertor solutions should be developed to manage the edge radiative cooling, while advanced core turbulence control should be explored to optimize core density and confinement, thereby extending the operational space of future reactors.

### v3

@@ III.3.1 Visualizing the HFS Radiation Front
The evolution of this front proceeds in clear stages. Early in the discharge ($t=2800$ ms), before the density ramp, the radiation is concentrated at the inner divertor target, characteristic of a standard attached divertor [Fig. 7(a)]. As plasma density increases ($t=3800$ ms), an intense radiation zone emerges near the X-point, signaling the birth of the radiative front [Fig. 7(b)]. This event coincides with the onset of the divergent turbulence seen in Fig. 6. With a further increase in density ($t=4800$ ms), this front detaches from the divertor and propagates upward along the HFS separatrix [Fig. 7(c)]. This movement is strongly correlated with the localized edge cooling and turbulence saturation. Finally ($t=5800$ ms), the front moves further inboard and its emissivity intensifies [Fig. 7(d)].
This sequence provides direct visual evidence of a developing radiative instability on the HFS. This dynamic structure appears to be the primary driver of the edge density limit. Notably, the eventual discharge termination was caused by a subsequent drop in NBI power that triggered an $n=1$ locked mode and was not a direct consequence of the radiation front.
III.3.2 Radiative Cooling and the Onset of HFS Turbulence
The formation of the HFS radiation front actively cools the plasma and triggers a dramatic shift in the local turbulent state. Figure 8 provides quantitative evidence of this transition. As the HFS radiated power increases beyond a critical threshold, the edge plasma parameters and HFS turbulence changes fundamentally. The RMS amplitude of the line-integrated HFS density fluctuations ($\tilde{n}_{\mathrm{HFS}}$), measured by the CO2 interferometer passing through the radiation front, serves as a proxy for the instability’s strength.
Figure 8: Transition to a degraded confinement state driven by HFS radiated power. (a) The edge adiabaticity parameter at $\rho\sim 0.9$ drops below unity as radiated power exceeds a critical threshold. (b) This transition coincides with a surge in the RMS amplitude of HFS density fluctuations ($\tilde{n}_{e}$). Data points are colored by the $H_{98y2}$ energy confinement factor, showing a clear shift from a high-confinement (yellow) to a low-confinement (purple) state. The plot is in log-log scale. 20 ms moving average is applied to fluctuation data.
The effect of edge cooling can be characterized by the dimensionless adiabaticity parameter [37], $\alpha_{\mathrm{adia}}=k^{2}_{\parallel}v^{2}_{te}/\omega\nu_{ei}\propto\frac{T_{e}^{5/2}}{n_{e}}$, which represents the comparison between parallel electron motion and turbulent timescales. Here, $k_{\parallel}$ is the parallel wavenumber of turbulence, $v_{te}$ is the electron thermal velocity, $\omega$ is the turbulence frequency, and $\nu_{ei}$ is the electron-ion collision frequency. When $\alpha_{\mathrm{adia}}<1$, the plasma enters a non-adiabatic regime where drift-wave instabilities are expected to grow. As shown in Fig. 8(a), the onset of the strong HFS radiation drives $\alpha_{\mathrm{adia}}$ to plummet below 1.
This transition into a non-adiabatic state coincides with a surge in HFS turbulence. Figure 8(b) shows that the RMS amplitude of HFS density fluctuations ($\tilde{n}_{e}$) increases sharply as the radiated power crosses the same threshold. The entire event is correlated with a degradation in plasma performance, as indicated by the drop in the $H_{98y2}$ confinement factor (color scale). This finding demonstrates that the HFS radiation front alters the local plasma state, unleashing the turbulence that may ultimately limit the edge density.
III.3.3 Clamping of the Edge Density
The preceding evidence indicates that the HFS radiative instability is the cause of the edge density saturation. To test this hypothesis quantitatively, we correlate the local edge density with the instability’s strength, again using the RMS amplitude of the line-integrated HFS density fluctuations ($\tilde{n}_{\mathrm{HFS}}$) as a proxy.
Figure 9 plots the edge density (at $\rho=0.95$) against $\tilde{n}_{\mathrm{HFS}}$ on a log-log scale. The data, compiled from multiple discharges, reveals a dramatic bifurcation in the plasma behavior, defined by two distinct operational regimes that are distinguished by the energy confinement factor, $H_{98y2}$.
@@ III.3.3 Clamping of the Edge Density
At low turbulence amplitudes, the plasma is in a high-confinement state (high $H_{98y2}$, yellow points). In this phase, the edge density and turbulence grow together, following a nearly linear relationship ($n_{e,\mathrm{edge}}\propto\tilde{n}_{\mathrm{HFS}}^{1.0}$). This represents a state where turbulence-driven transport has not yet begun to limit the density.
This linear relationship breaks down abruptly once the turbulence level surpasses a critical threshold, at which point the system transitions to a degraded-confinement state (low $H_{98y2}$, purple points). While the instability amplitude continues to grow by nearly an order of magnitude, the edge density saturates almost completely, exhibiting a nearly flat scaling ($n_{e,\mathrm{edge}}\propto\tilde{n}_{\mathrm{HFS}}^{0.1}$).
This bifurcation provides strong evidence that the edge density limit is not a gradual process, but a sharp transition triggered by the radiative instability. Once initiated, the instability clamps the edge density at a value consistently below the Greenwald limit, leading to a significant increase in HFS turbulence. The apparent overall relationship across both regimes ($n_{e,\mathrm{edge}}\propto\tilde{n}_{\mathrm{HFS}}^{0.23}$) also masks the sharpness of this critical transition.
III.3.4 Breakdown of Zonal Flow Regulation
@@
The onset of the HFS radiative instability has global consequences, leading to a failure of turbulence self-regulation across the plasma edge. We use BES measurements on the low-field side to observe this breakdown in detail (Fig. 10), revealing the microphysical mechanism responsible for enhanced transport.
The trigger for this breakdown is the transition to the non-adiabatic regime. As the edge cools with the onset of the radiation front ($t\approx 3700$ ms), the dimensionless adiabaticity parameter, $\alpha_{\mathrm{adia}}$, drops below unity [Fig. 10(a)]. This transition is known to fundamentally alter drift-wave turbulence, weakening the mechanisms that generate stabilizing zonal flows [37, 13].
While the fluctuation amplitude on the LFS saturates (as seen in Fig. 6), the underlying dynamics show a flattening of the Reynolds stress profile [$\langle\tilde{v}_{\theta}\tilde{v}_{r}\rangle$, Fig. 10(d)]. This collapse of the primary nonlinear drive for zonal flows (i.e., the Reynolds force, $\mathcal{F}_{\text{Re}}=-\partial_{r}\langle\tilde{v}_{\theta}\tilde{v}_{r}\rangle$) signifies a breakdown in the zonal flow regulation mechanism. Concurrently, the mean poloidal phase velocity of turbulence [$\langle\tilde{v}_{\theta}\rangle$, Fig. 10(c)] also collapses.
@@
This breakdown in zonal flow regulation directly drives particle loss. The outward turbulent particle flux [$\langle\tilde{n}\tilde{v}_{r}\rangle$, Fig. 10(e)] grows significantly, while the turbulent spreading flux [$\langle\tilde{n}^{2}\tilde{v}_{r}\rangle$, Fig. 10(f)] indicates that turbulence intensity is propagating inward from the separatrix. These measurements suggest a compelling narrative: edge cooling initiates a transition to the non-adiabatic turbulent regime, which disables zonal flow regulation and unleashes enhanced transport.
III.4 The Core Density Limit: A Transition to Avalanche-like Turbulence
While the edge density saturates below the Greenwald limit, the core density continues to rise, often reaching much higher values than $n_{\text{G}}$. This decoupling implies that a separate, core-localized mechanism must govern the core density profile. Here, we present evidence that this mechanism is not a simple enhancement of diffusive transport, but a fundamental transition in its character to intermittent, avalanche-like behavior. This transition is enabled by the collapse of the stabilizing $E_{r}\times B$ shear, as shown previously in Fig. 5.
III.4.1 Statistical Signatures of Avalanche-like Transport,
The statistical properties of the core density fluctuations, measured by the RIP diagnostic at the midplane, reveal three distinct signatures of this transition.
First, the probability density function (PDF) of the fluctuation amplitudes evolves from Gaussian to heavy-tailed. Early in the discharge the PDF is Gaussian [yellow in Fig. 11(a)], characteristic of diffusive transport. As the density limit is approached, however, the PDF develops pronounced, non-Gaussian tails [purple in Fig. 11(a)]. This signifies a shift to a transport regime dominated by large-amplitude, intermittent events. The growing intermittency is quantified by the kurtosis of the PDF, which rises from the Gaussian value of 3 to about 6 [Fig. 11(b)].
@@ III.4.1 Statistical Signatures of Avalanche-like Transport,
Second, the frequency spectrum of the turbulence develops a $1/f$-type power-law scaling. As the plasma evolves with rising density, a distinctive power-law scaling emerges at a wide range of frequencies [Fig. 12(a)]. Such a spectrum is a classic signature of self-organized criticality (SOC) and avalanche dynamics, where transport occurs via cascading events spanning a wide range of scales [38, 39, 40].
Third, the transport acquires long-range temporal correlations, quantified by the Hurst exponent ($H$), a measure of the persistence in a time series [Fig. 12(b)]. The Hurst exponent rises from $H\approx 0.5$, the value for uncorrelated, random-walk events (diffusive transport), to $H\approx 0.8$ as the density limit is approached. A value of $H>0.5$ indicates persistent, correlated dynamics, another key feature of avalanche-like transport. This rise in $H$ is strongly correlated with a substantial decrease in the energy confinement factor $H_{98y2}$, linking the emergence of this non-diffusive transport regime to the degradation in plasma performance.
Collectively, these three independent signatures—heavy-tailed PDFs, a $1/f$ power spectrum, and an elevated Hurst exponent—provide a consistent and compelling picture: core transport undergoes a fundamental transition from a diffusive to an intermittent, avalanche-like regime, which may be the mechanism for the core density limit.
@@ III.4.2 Core Density Saturation by Avalanche-like Turbulence
The emergence of avalanche-like dynamics provides a clear hypothesis for the core density limit: the large, intermittent transport events ultimately clamps the density profile. We test this hypothesis by correlating the core density (at $\rho\approx 0.7$) with the RMS amplitude of the midplane turbulence measured by RIP.
@@
As shown in Fig. 13, data compiled from multiple discharges reveal a clear bifurcation between two transport regimes, distinguished by their confinement properties ($H_{98y2}$).
In the low-turbulence, high-confinement regime ($H_{98y2}>1.0$, yellow points), which corresponds to the diffusive state, the core density rises robustly with increasing turbulence amplitude, following a strong power-law scaling of $n_{e,\mathrm{core}}\propto\tilde{n}_{\mathrm{rms}}^{0.49}$. This phase represents a state where fueling remains effective despite the presence of turbulence.
This behavior changes dramatically as the transport transitions to the high-turbulence, low-confinement regime ($H_{98y2}<1.0$, purple points), which corresponds to the avalanche-like state. Here, the scaling weakens significantly, flattening to $n_{e,\mathrm{core}}\propto\tilde{n}_{\mathrm{rms}}^{0.17}$. This transition to a much weaker scaling demonstrates that the core fueling has become highly inefficient. The avalanche-like fluctuations effectively prevent any significant increases in density, leading to a saturation of the core density profile. This saturation occurs at densities far exceeding the Greenwald limit, approaching $2n_{\mathrm{G}}$.
The weak overall scaling of density with turbulence ($n_{\mathrm{core}}\propto\tilde{n}_{\mathrm{rms}}^{0.26}$) is consistent with the weak macroscopic power scaling of the core density limit ($n_{e,\mathrm{core}}\propto P_{\mathrm{SOL}}^{0.27\pm 0.03}$) found previously. This alignment confirms that the core density limit is governed not by simple power balance, but by a fundamental transition in the turbulent transport itself, from diffusive to avalanche-like, which saturates the density profile.
@@ IV Conclusion and Discussion
This study re-examined the underlying physics of the density limit using negative triangularity plasmas on the DIII-D tokamak. Our findings reveal that the density limit, at least in NT plasmas without an H-mode edge, is not a singular disruptive boundary, but a multifaceted saturation phenomenon governed by distinct physical mechanisms: an edge density limit triggered by a radiative instability, and a core density limit driven by avalanche-like turbulent transport. Moreover, this study demonstrate sustained, non-disruptive operations at densities up to 1.8 times the Greenwald limit with substantial auxiliary heating.
At the plasma edge, the density limit manifests as an abrupt saturation event triggered by a radiative instability on the high-field side. While geometrically similar to a disruptive MARFE, this radiative instability acts as a quasi-stable front that clamps the edge density at a level below the Greenwald limit, without terminating the discharge.
In contrast, the core density is not directly constrained by the edge cooling and continues to rise well beyond the Greenwald limit. Its saturation is instead governed by a fundamental shift in the nature of core turbulence. As the density increases, core transport transitions from a diffusive state to one characterized by intermittent, avalanche-like dynamics. This is supported by signatures including heavy-tailed PDFs, $1/f$-type power spectra, and elevated Hurst exponents. This persistent, non-diffusive transport regime prevents further increases in the core density.
@@
These findings also present a potential strategy for achieving high Greenwald fraction operation in future tokamaks, particularly those utilizing negative triangularity configurations. Success hinges on addressing the distinct physical mechanisms governing the density limit in the core and edge. The edge radiative instability can be mitigated through advanced divertor designs and impurity control aimed at managing boundary heat fluxes. Concurrently, advanced techniques to suppress core turbulence, such as strong rotational shear flows and Shafranov shift, can be developed to optimize density profiles and confinement. By pursuing these parallel control strategies, the operational space of future reactors can be extended into stable, high-density regimes well beyond conventional scaling.